OpenStudy (anonymous):

what constant must be added to x^4+8^2+ to make it perfect square trinomial ?

OpenStudy (whpalmer4):

Is that supposed to be \[x^4+8x^2\]? a perfect square is of the form \[(u+a)^2 = u^2 + 2au + a^2\]If we compare that with yours: \[x^4+8x^2\]\[u^2+2au + a^2\] \[x^4=u^2\]\[x^2=u\] \[8x^2 = 2au\]\[8=2a\]\[4=a\] So we need to add what to our expression to make it a perfect square?

OpenStudy (anonymous):

[x^2 + 4]^2 = x^4 + 8x^2 + 16 Answer: Add 16